Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

five ordinary quarters and a fake quarter with two heads are places in a hat. One quarter is selected at random and tossed twice. If the outcome is "HH", what is the probability that the fake quarter was selected?

OpenStudy (anonymous):

1/6 I'd say

OpenStudy (anonymous):

Ok, You have observed HH. The probability of getting that with a fair coin is 1/4, and you have a 4/5 chance of getting that. The probability of getting HH with your fake coin is 1 (has to happen), but you have a 1/5 chance of getting this. So there is 1/4*4/5 + 1/5 = 2/5 chance that you get HH for a random coin in the bag. But you got HH, so you treat that as if it is 100%, and your answer is P(fake HH)/P(total HH)=1/5 / 2/5 = 1/2 or 0.5

OpenStudy (anonymous):

functional, how do you have a 1/5 chance of getting the fake coin if there are 6 coins in the bag?

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!